Two painters, Jane and Jesse, working together paint a certain room in 3 hours. If it takes Jesse, working alone, 8 hours to paint the room, how long would it take Jane to paint the same room all by herself?
Answer by Boreal(15235)

Work these as fractions of room painted in 1 hour.
Jesse (1/8)
J and J (1/3)
Therefore
(1/8)+ x=(1/3)
Multiply all by 24 to clear fractions
3+24x=8
24x=5
x=5/24
Jane paints 5/24th of the room in 1 hour.
It takes her 24/5 hours or 4.8 hours to paint the room.